1. Cabell Products is a division of a major...

Cabell Products is a division of a major corporation. Last year the division had total sales of $27,220,000, net operating income of $2,874,320, and average operating assets of $7,900,000. The company's minimum required rate of return is 12%.

The division's turnover is closest to:

Multiple Choice

  • 9.47

  • 3.45

  • 0.36

  • 3.03

2.

Cabell Products is a division of a...

Cabell Products is a division of a major corporation. Last year the division had total sales of $12,270,000, net operating income of $834,360, and average operating assets of $3,190,200. The company's minimum required rate of return is 12%.

The division's return on investment (ROI) is closest to:

Multiple Choice

  • 6.8%

  • 26.2%

  • 56.7%

  • 26.0%

3.

The Consumer Products Division of...

The Consumer Products Division of Goich Corporation had average operating assets of $950,000 and net operating income of $96,600 in May. The minimum required rate of return for performance evaluation purposes is 10%.

What was the Consumer Products Division's residual income in May?

Multiple Choice

  • $(1,600)

  • $9,660

  • $1,600

  • $(9,660)

4.

Wallen Corporation is considering eliminating ...

Wallen Corporation is considering eliminating a department that has an annual contribution margin of $80,000 and $160,000 in annual fixed costs. Of the fixed costs, $90,000 cannot be avoided. The annual financial advantage (disadvantage) for the company of eliminating this department would be:

Multiple Choice

  • $10,000

  • ($10,000)

  • $80,000

  • ($80,000)

5.

The following information relates ...

The following information relates to next year's projected operating results of the Children's Division of Grunge Clothing Corporation:

Contribution margin $ 200,000
Fixed expenses 500,000
Net operating loss $ (300,000 )

If the Children's Division is eliminated, $170,000 of the above fixed expenses could be avoided. The annual financial advantage (disadvantage) for the company of eliminating this division should be:

Multiple Choice

  • ($300,000)

  • $30,000

  • ($30,000)

  • $300,000
Answer #1

1. Turnover = Sales /Average operating assets

= 27,220,000/7,900,000

= 3.45

Option B

2.

Margin = Net operating income /sales

= 834,360/12,270,000 = 6.8%

Turnover = Sales /Average operating assets

=12,270,000/3,190,200 = 3.85

ROI = margin *turnover = 6.8*3.85 =26.2%

Option B

3.

Residual income = Net operating income - required return

= 96,600 - (950,000*10%)

= 1,600

Option C